OpenStudy (anonymous):

Two panes are headed towards eachother, an are 440 iles apart. If one is going 240 miles an hour and the other 360 mph how far will they be four minutes before meeting?

OpenStudy (anonymous):

40 miles. their total speed is 600 miles per hour. the total distance is 440 so the total time is \[\frac{440}{600}=\frac{11}{15}\] that is eleven-fifteenth of an hour. one fifteenth of an hour is 4 minutes, so eleven fifteenths is 44 minutes. 4 minutes before they crash into each other is therefore 40 minutes. if they travel at 600 mph for 40 minutes they go \[600\times \frac{2}{3}=400\] because 40 minutes is two-thirds of an hour. therefore they are 40 miles apart.